OpenStudy (anonymous):

what is the equation of the circle with center (0, 0) that passes through the point (-5, -5)?

hartnn (hartnn):

Equation of circle passing through centre is of the form \(\huge x^2+y^2=r^2\) now just plug in x= -5 y=-5 and find r :)

OpenStudy (anonymous):

I know how to do it but I don't know if my answer is right =/

hartnn (hartnn):

what answer you got ?

OpenStudy (anonymous):

x2 + y2 = 50

hartnn (hartnn):

thats absolutely correct! :)

OpenStudy (anonymous):

thank you! can you help me with few more?

hartnn (hartnn):

sure :)

OpenStudy (haseeb96):

good job @hartnn

OpenStudy (haseeb96):

and @Jessica_sunshinexx u too...

OpenStudy (anonymous):

what are the center and radius of the circle with equation (x-1)^2 + (y+4)^2 =4?

hartnn (hartnn):

the standard equation is \(\huge (x-h)^2 +(y-k)^2 =r^2\) where, (h,k) is the center of the circle and r= radius of circle

hartnn (hartnn):

you just need to compare the equation in the question to find the values of h,k and r^2

hartnn (hartnn):

did u get what to do ?

OpenStudy (anonymous):

center 1 , -4 r + 2 is that the answer? @hartnn

hartnn (hartnn):

absolutely correct! :)